# BiCo-Fusion: Bidirectional Complementary LiDAR-Camera Fusion for Semantic- and Spatial-Aware 3D Object Detection

### v1

@@ Abstract
3D object detection is an important task that has been widely applied in autonomous driving. Recently, fusing multi-modal inputs, i.e., LiDAR and camera data, to perform this task has become a new trend. Existing methods, however, either ignore the sparsity of Lidar features or fail to preserve the original spatial structure of LiDAR and the semantic density of camera features simultaneously due to the modality gap. To address issues, this letter proposes a novel bidirectional complementary Lidar-camera fusion framework, called BiCo-Fusion that can achieve robust semantic- and spatial-aware 3D object detection. The key insight is to mutually fuse the multi-modal features to enhance the semantics of LiDAR features and the spatial awareness of the camera features and adaptatively select features from both modalities to build a unified 3D representation. Specifically, we introduce Pre-Fusion consisting of a Voxel Enhancement Module (VEM) to enhance the semantics of voxel features from 2D camera features and Image Enhancement Module (IEM) to enhance the spatial characteristics of camera features from 3D voxel features. Both VEM and IEM are bidirectionally updated to effectively reduce the modality gap. We then introduce Unified Fusion to adaptively weight to select features from the enchanted Lidar and camera features to build a unified 3D representation. Extensive experiments demonstrate the superiority of our BiCo-Fusion against the prior arts. Project page: https://t-ys.github.io/BiCo-Fusion/.
@@ I INTRODUCTION
3D object detection [1, 2, 3] is a critical and challenging task that aims to localize and classify objects in the 3D space, which has been widely applied in applications such as robotics and autonomous driving. Early works often resorted to one single sensor as inputs, such as LiDARs or RGB cameras. While approaches based on these two sensors have yielded significant results, yet they both have their own unsolvable problems. LiDAR-based methods [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have difficulty detecting long-distance or small objects due to the sparsity of point cloud and the lack of semantics, while camera-based methods [19, 20, 21, 22, 23, 24, 25, 26] perform even worse, struggling in localization due to the lack of accurate spatial and depth information, especially when the objects are occluded.
Recently, a new trend to handle these problems is to fuse two modalities, aiming to utilize more complementary information from both sparse point clouds and dense camera data. In the context of this trend, existing methods to deal with the fusion of two heterogeneous modalities can be divided into two main categories: i) 2D-plane fusion (Fig. 1(a)) and ii) 3D-space fusion (Fig. 1(b)). 2D-plane fusion methods [27, 28, 29, 30, 31, 32, 2] perform LiDAR-camera fusion by projecting point clouds to the image plane and retrieving the nearby 2D features, which are used to decorate the 3D representations for compensating the semantic capabilities. However, such a projection ignores the sparsity of LiDAR data, and inevitably wastes semantically rich 2D features [33].
3D-space fusion methods, on the other hand, unify two modalities in the same space [34, 1, 3], either in the voxel space [3] or in bird’s eye view (BEV) space [1]. During the fusion in a unified space, the sparsity of geometry is well resolved by the dense features from the camera, making the performance significantly stronger than that of 2D-plane fusion.
@@
Figure 1: Comparison of our framework with previous LiDAR-camera fusion methods. (a) 2D-plane fusion projects point clouds to the image plane and retrieves the corresponding 2D features which are used to augment original points. (b) 3D-space fusion firstly estimates the depth of images and then lifts them by view transform to a unified space, where the two modalities are fused. (c) Our BiCo-Fusion proposes a bidirectional complementary fusion to reduce the modality gap and achieve a smoother process. In the initial Pre-Fusion stage, we introduce VEM and IEM to apply specific enhancements to each modality which provide semantics for LiDAR and spatial capability for camera, respectively. During Unified Fusion, the enhanced features are unified in a 3D space by view transformation for further adaptive-weighted-based fusion.
Despite the impressive improvements, the modality gap makes it fail to preserve the original strengths of these two modalities simultaneously. Specifically, although it is expected to fuse the strong representations from two modalities, i.e., the strong spatial capability of LiDAR and the dominant semantic density of the camera. It can not be neglected that these two modalities also have their drawbacks (i.e., LiDAR lacks semantics and camera struggles in spatial perception), which are also fused when building a unified 3D representation. Such a ‘negative effect’ renders the fused features often imperfect. In other words, it weakens the impact of the strong aspects of the original modality, leading to a substantial loss of modality-specific strengths.
@@
In this letter, we propose BiCo-Fusion, a novel multi-modal 3D object detection framework to tackle the above problems. BiCo-Fusion employs a bidirectional complementary strategy to mutually fuse the multi-modal features to enhance the semantics of LiDAR features and the spatial awareness of the camera features and adaptatively select features from both modalities to build a unified 3D representation, as shown in Fig. 1(c). This subtly ensures semantic and spatial awareness and reduces the modality gap, thereby facilitating a smoother and more efficient fusion. Firstly, the extracted features from LiDAR and the camera are input into the initial stage of fusion, referred to as Pre-Fusion (Sec. III-B). Here, two modules play crucial roles: the Voxel Enhancement Module (VEM) projects non-empty voxels onto the image plane and extracts the surrounding camera features to enhance their semantics, which is optimized based on our distance-prior weighting scheme. The Image Enhancement Module (IEM) fuses the real depth from the LiDAR with camera features to enhance their spatial awareness. Through Pre-Fusion, the drawbacks of both modalities undergo specific enhancements, avoiding the negative effect, thus reducing the modality gap. Next, we lift the enhanced camera features to a unified voxel space by view transformation [19], which is fused with the enhanced LiDAR voxel features. We then introduce an adaptive weighting scheme to perform Unified Fusion to dynamically select the 3D features that possess both spatial-semantic awareness (Sec. III-C). Overall, these components work seamlessly to guarantee a more robust 3D representation for the detection.
@@
Extensive experiments on the nuScenes [35] benchmark demonstrate that our BiCo-Fusion approach achieves state-of-the-art (SoTA) performance, with 72.4% mAP and 74.5% NDS on the test set. In summary, our contributions are three-fold: (I) We propose a novel multi-modal fusion framework for 3D object detection task, employing bidirectional complementary LiDAR-camera fusion to preserve the spatial structure from LiDAR and the semantics from camera, which reduces the modality gap, allowing for a more comprehensive 3D representation; (II) We propose Pre-Fusion with VEM for enhancing the semantics of voxels and IEM for enhancing the spatial awareness of camera data, and Unified Fusion for adaptively unifying representations in the 3D voxel space; (III) We achieve the SOTA 3D detection performance on the challenging nuScenes dataset.
@@ II RELATED WORK
LiDAR-based 3D Object Detection. LiDAR point clouds are inherently suitable for 3D object detection, in that they can provide accurate 3D spatial information. However, the point cloud data is sparse, disordered, and redundant, which results in point cloud data that cannot be directly fed into a CNN network for feature extraction. PointNet [36] is the first backbone capable of extracting point cloud features, After that, a great deal of subsequent works, e.g., [17, 8, 16] have been proposed. Later, to obtain more regular point cloud data, researchers choose to rasterize point cloud data into discrete grid representation, such as voxels [4, 13, 15] and pillars [5]. These approaches enable direct feature extraction with convolutional networks. However, they cannot break the limitation of using a single modality and are prone to false detections due to missing semantics in special cases, such as small and distant objects. In contrast, our method uses both LiDAR and camera data and mutually fuses the multi-modal features to enhance the semantics of LiDAR features and the spatial awareness of the camera features.
LiDAR-Camera Fusion-based 3D Object Detection. This has become a new trend, dedicated to maximizing the use of both worlds. Early works typically use a projection strategy to retrieve image features for decorating point clouds. PointPainting [27] is a pioneering work to enhance semantic capabilities by projecting semantic segmentation labels of images into the point cloud space. While MV3D [28] projects 3D proposals into 2D plane for connecting two modalities. TransFusion [2] uses the attention mechanism to selectively fuse relevant image features to the point cloud. Thanks to lift-splat-shoot (LSS) [19], recent research efforts are mostly based on it to transform camera features into the same unified space with LiDAR. BEVFusion [1] provides a new pipeline that unifies the two modalities in birds’ eye view (BEV) space. While UVTR [3] uses the two modalities by unifying them in voxel space to avoid the dislocation effect of height collapse. However, these methods fail to preserve the original spatial structure of LiDAR and the semantic density of camera features simultaneously due to the modality gap. By contrast, our BiCo-Fusion proposes a bidirectional complementary strategy to mutually fuse the multi-modal features to enhance the semantics of LiDAR features and the spatial awareness of the camera features, which are adaptatively weighted to build a unified 3D representation.
@@ III-A Framework Overview
Then, the extracted LiDAR voxel features $F_{L}$ and camera features $F_{C}$ are interacted with and fused using our designed method. Our multi-modal fusion space contains two fusions, in the Pre-Fusion phase, our Voxel Enhancement Module (VEM) and Image Enhancement Module (IEM) modules perform specific bidirectional complementary enhancement for both modalities to improve the semantic capability of voxels and the spatial capability of images, which effectively avoids the negative effects caused by the weak representations, and also prepares more comprehensive features for the next step. Afterward, in the Unified Fusion phase (U-Fusion), the enhanced spatial-aware camera features $F_{SpC}$ then be lifted into the 3D voxel space by view transformation [19, 1, 3], which are further fused with the semantic-aware LiDAR voxel features $F_{SeL}$ in an adaptive way.
Finally, following the workflow of the voxel-based approaches [4, 6, 1], we collapse the height of the fused voxel features $F_{f}$ in order to transform them into the BEV space, then the BEV features $F_{B}$ go through the convolutional-based BEV encoder and the detection head for obtaining the final 3D detection results. Following UVTR [3], the Hungarian algorithm is applied to match the predictions and the ground truth during training. Meanwhile, Focal loss [38] and $L1$ loss are used for the classification and 3D bounding box regression, respectively.
@@ III-B Pre-Fusion
The Pre-Fusion consists of two key modules: the VEM and the IEM. These two modules work in a bidirectional complementary manner to enhance the semantics of voxels and the spatial characteristics of images. It mitigates the weak representations of LiDAR and camera. Meanwhile, the strong representations of them are preserved, allowing for a more comprehensive 3D representation.
Figure 4: Image Enhancement Module (IEM). It first projects raw LiDAR points to the image coordinates. Then, we use depth completion to generate a dense depth feature map, whose size is the same as the camera features. Finally, the dense depth feature map is concatenated with the camera features, followed by a convolution layer to reduce the feature dimension.
@@ III-B1 Voxel Enhancement Module
Despite having strong spatial localization capabilities, the LiDAR voxel features $F_{L}$ obtained from the LiDAR branch lack important semantic information. This drawback can propagate through the Unified Fusion process. To overcome this issue and provide more comprehensive features for the Unified Fusion, we design the VEM, as shown in Fig. 3.
@@
$F_{weighted}=\operatorname{Softmax}(L_{nearest})\times F_{nearest}\text{,}$ (1)
@@
$F_{SeL}=\operatorname{ReLU}(\operatorname{Linear}(F_{weighted}))+F_{L}\text{.}$ (2)
@@ III-B2 Image Enhancement Module
Camera features lack the perception ability of 3D spatial information, which is crucial for 3D object detection tasks. To enhance them from this view, we propose IEM, as shown in Fig. 4. We first project the point cloud onto the image plane to obtain a sparse depth map $D_{sparse}$. Inspired by [39], we then use depth completion [40] followed by feature extraction to obtain a dense depth feature map $D_{dense}\in\mathbb{R}^{H\times W\times C_{depth}}$. Finally, we concatenate the dense depth feature map with the camera features $F_{C}\in\mathbb{R}^{H\times W\times C_{2D}}$ and use a convolutional layer to fuse them, getting the spatial-aware camera features $F_{SpC}$:
$F_{SpC}=\operatorname{Conv}(\operatorname{Concat}(F_{C},D_{dense})\text{.}$ (3)
Due to the geometric capability of the spatially-aware camera features, lifting them into 3D space ensures a more accurate process, providing a solid foundation for the subsequent Unified Fusion.
@@ III-C Unified Fusion
After obtaining the semantic-aware LiDAR voxel features and spatial-aware camera features, which have undergone specific bidirectional complementarity during Pre-Fusion, we further fuse them in the U-Fusion stage to ensure the integrity of strong representations, for both semantic and spatial levels.
Previous work [1, 34] unifies the two modalities in the BEV space by view transform. While this improved efficiency, it has an inherent limitation that prevents the fine-grained fusion which is vital for better performance. Therefore, after elevating the camera features to the voxel representation, we preserve its state, obtaining $\hat{F}_{SpC}$.

### v3

@@ Abstract
3D object detection is an important task that has been widely applied in autonomous driving. To perform this task, a new trend is to fuse multi-modal inputs, i.e., LiDAR and camera. Under such a trend, recent methods fuse these two modalities by unifying them in the same 3D space. However, during direct fusion in a unified space, the drawbacks of both modalities (LiDAR features struggle with detailed semantic information and the camera lacks accurate 3D spatial information) are also preserved, diluting semantic and spatial awareness of the final unified representation. To address the issue, this letter proposes a novel bidirectional complementary LiDAR-camera fusion framework, called BiCo-Fusion that can achieve robust semantic- and spatial-aware 3D object detection. The key insight is to fuse LiDAR and camera features in a bidirectional complementary way to enhance the semantic awareness of the LiDAR and the 3D spatial awareness of the camera. The enhanced features from both modalities are then adaptively fused to build a semantic- and spatial-aware unified representation. Specifically, we introduce Pre-Fusion consisting of a Voxel Enhancement Module (VEM) to enhance the semantic awareness of voxel features from 2D camera features and Image Enhancement Module (IEM) to enhance the 3D spatial awareness of camera features from 3D voxel features. We then introduce Unified Fusion (U-Fusion) to adaptively fuse the enhanced features from the last stage to build a unified representation. Extensive experiments demonstrate the superiority of our BiCo-Fusion against the prior arts. Project page: https://t-ys.github.io/BiCo-Fusion/.
@@ I INTRODUCTION
3D object detection [1, 2, 3] is a critical and challenging task that aims to localize and classify objects in the 3D space, which has been widely applied in applications such as robotics and autonomous driving. Early works often resorted to one single sensor as inputs, such as LiDARs or RGB cameras. While approaches based on these two sensors have yielded significant results, they yet need to address and overcome several inherent limitations. LiDAR-based methods [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have difficulty classifying small objects due to poor performance on detailed semantic information, while camera-based methods [19, 20, 21, 22, 23, 24, 25, 26] perform even worse, struggling with localization due to the lack of accurate spatial and depth information, especially when the objects are occluded [27].
A new trend to handle these problems is to fuse two modalities, aiming to utilize more complementary information from both sparse point clouds and dense camera data. In the context of this trend, existing methods to deal with the fusion of two heterogeneous modalities can be divided into two main categories: i) 2D-plane fusion (Fig. 1(a)) and ii) 3D-space fusion (Fig. 1(b)). Early 2D-plane fusion methods [28, 29, 30, 31, 32, 33, 2] perform LiDAR-camera fusion by projecting point clouds or proposals generated from them to the image plane and retrieving the corresponding 2D camera features, which are used to decorate the 3D representations in order to compensate for the semantic ability. However, the camera features are much denser, such a projection only considers fusing sparse point clouds or proposals with part of them, which wastes semantically rich 2D features [34].
While recent 3D-space fusion methods unify two modalities in the same 3D space [35, 1, 3], either in voxel [3] or in bird’s eye view (BEV) space [1], and fuse them to obtain a unified representation. This significantly solves the problem of 2D-plane fusion as all dense camera features can be transformed into 3D space to participate in the fusion with LiDAR.
@@
Figure 1: Comparison of our framework with previous LiDAR-camera fusion methods. (a) 2D-plane fusion projects point clouds or proposals generated from them to the image plane and retrieves 2D features for fusion. (b) 3D-space fusion lifts camera features by view transformation to a unified space with LiDAR features, where the two modalities are fused. (c) Our BiCo-Fusion proposes Pre-Fusion with a bidirectional complementary strategy to compensate for the drawbacks of both modalities, including VEM and IEM which enhance semantic awareness for LiDAR and 3D spatial awareness for the camera, respectively. During Unified Fusion, the enhanced features are unified in a 3D space by view transformation for further adaptive-weighting-based fusion.
However, it cannot be ignored that both modalities have drawbacks. Specifically, LiDAR features struggle with detailed semantic information while the camera falls short of accurate 3D spatial information. These drawbacks are also inevitably preserved during direct fusion in a unified space. As a result, this dilutes semantic and spatial awareness of the final unified representation, which are modality-specific strengths inherited from the camera and LiDAR, respectively.
@@
In this letter, we propose BiCo-Fusion, a novel multi-modal 3D object detection framework to tackle the above problem. BiCo-Fusion employs a bidirectional complementary strategy to fuse LiDAR and camera features to enhance the semantic awareness of the LiDAR and the 3D spatial awareness of the camera. It then adaptively fuses the enhanced features from both modalities to construct a semantic- and spatial-aware unified representation, as shown in Fig. 1(c). Firstly, the extracted features from LiDAR and the camera are input into the initial stage of fusion, referred to as Pre-Fusion (Sec. III-B). Here, two modules work together in a bidirectional complementary way to play crucial roles: the Voxel Enhancement Module (VEM) projects non-empty voxels onto the image plane and extracts the surrounding camera features to enhance their semantic awareness, which is optimized based on our distance-prior weighting scheme. The Image Enhancement Module (IEM) fuses the real depth from the LiDAR with camera features to enhance their 3D spatial awareness. Through Pre-Fusion, the drawbacks of both modalities undergo our specific enhancements, preparing more comprehensive LiDAR and camera features for the subsequent process. Next, we unify the enhanced camera features to the same 3D space as LiDAR by view transformation [19]. Finally, we introduce an adaptive weighting scheme to perform Unified Fusion (U-Fusion), which dynamically fuses the enhanced features from both modalities to build a unified representation (Sec. III-C). Overall, these components work seamlessly to guarantee a robust semantic- and spatial-aware 3D representation for the detection.
@@
Extensive experiments on the nuScenes [36] benchmark demonstrate that our BiCo-Fusion approach achieves state-of-the-art (SOTA) performance, with 72.4% mAP and 74.5% NDS on the test set. In summary, our contributions are three-fold: (I) We propose a novel BiCo-Fusion framework that enables a semantic- and spatial-aware unified representation for 3D object detection; (II) We propose Pre-Fusion works in a bidirectional complementary way, including VEM for enhancing the semantic awareness of LiDAR voxels and IEM for enhancing the 3D spatial awareness of camera features, and Unified Fusion for adaptively unifying representations in the 3D space; (III) We achieve the SOTA 3D detection performance on the challenging nuScenes dataset.
@@ II RELATED WORK
LiDAR-based 3D Object Detection. LiDAR point clouds are inherently suitable for 3D object detection, in that they can provide accurate 3D spatial information. However, the point cloud data is sparse and unordered, which results in point cloud data that cannot be directly fed into a CNN network for feature extraction. PointNet [37] is the first backbone capable of extracting point cloud features, inspiring a great deal of subsequent works [17, 8, 16]. Later, to obtain more regular point cloud data, researchers choose to rasterize point cloud data into discrete grid representation, such as voxels [4, 13, 15] and pillars [5], enabling direct feature extraction with convolutional networks. However, they cannot break the limitation of using a single modality and are prone to false detections due to missing semantics in special cases, such as small and distant objects.
LiDAR-Camera Fusion-based 3D Object Detection. This has become a new trend, aiming to maximize the use of both worlds. Early 2D-plane fusion methods (Fig. 1(a)) [28, 32, 29, 2, 31] typically use a projection strategy to retrieve corresponding camera features for decorating point clouds or proposals that are generated from them. PointPainting [28] is a pioneering work that performs fusion at the point level, which directly projects raw point clouds into the image plane for fusion. While MV3D [29] chooses to fuse camera features with the projected proposals, denoted as proposal-level methods [2]. However, regardless of point-level or proposal-level methods, they only associate the sparse LiDAR point clouds or proposals with part of the dense camera features, wasting semantically rich information in them.
Recently, 3D-space fusion methods (Fig. 1(b)) [34, 1, 3, 38, 39, 35] try to construct a unified representation for both modalities by unifying them in a shared 3D space for fusion. This paradigm naturally mitigates the problem of 2D-plane fusion as all dense camera features can be transformed into the shared space to engage in the fusion. UVTR [3] fuses the LiDAR and camera features in a unified voxel space. While BEVFusion [1] generates a unified representation by unifying both modalities in the birds’ eye view (BEV) space to improve the computational efficiency, which provides a solid pipeline for recent works [34, 38, 39, 27, 40]. However, these methods ignore that during direct fusion in a unified space, the drawbacks of both modalities (LiDAR features struggle with detailed semantic information and the camera lacks accurate 3D spatial information) are also preserved. It dilutes semantic and spatial awareness of the final unified representation.
To this end, our BiCo-Fusion (Fig. 1(c)) proposes Pre-Fusion with a bidirectional complementary strategy, including Voxel Enhancement Module (VEM) to enhance the semantic awareness of LiDAR features and Image Enhancement Module (IEM) to enhance the 3D spatial awareness of camera features. This helps to ensure the comprehensiveness of the two modalities at the semantic and spatial levels, beneficial for building up a unified representation in our Unified Fusion (U-Fusion) stage.
@@ III-A Framework Overview
Then, the extracted LiDAR voxel features $F_{L}$ and camera features $F_{C}$ are interacted with and fused using our designed method. Our multi-modal fusion space contains two fusions, in the Pre-Fusion phase, our Voxel Enhancement Module (VEM) and Image Enhancement Module (IEM) perform enhancements for both modalities in a bidirectional complementary way to improve the semantic capability of voxels and the spatial capability of images, which effectively compensates for their drawbacks, and also prepares more comprehensive features for the next step. Afterward, in the Unified Fusion phase (U-Fusion), the enhanced spatial-aware camera features $F_{SpC}$ then be unified into the 3D voxel space by view transformation [19, 1, 3], which are further fused with the semantic-aware LiDAR voxel features $F_{SeL}$ adaptively to build a unified representation $F_{f}$.
Following the workflow of voxel-based approaches [4, 6, 1], we collapse the height of the fused voxel features $F_{f}$ to transform them into BEV space. Then, we follow [1] to introduce a convolution-based BEV encoder containing several convolutional layers to compensate for spatial misalignments between the LiDAR and camera caused by inaccurate depth during view transformation.
Following prior works [1, 35, 2, 34], we use the same detection head as in [2] that is built upon transformer architecture [42], which is first proposed in DETR [43]. The detection head consists of a transformer decoder and a feed-forward network. A set of embeddings (called object queries) [42] serves as the input for the decoder. During the decoder, the object queries are passed through a self attention module, then cross attention is conducted with our final fused BEV features $F_{B}$. The cross attention between object queries and the feature maps aggregates relevant context onto the object candidates, while the self attention between object queries reasons pairwise relations between different object candidates. After the decoder, the object queries containing rich instance information are then independently decoded into boxes and class labels by a feed-forward network (FFN).
Following UVTR [3], the Hungarian algorithm is applied to match the predictions and the ground truth during training. Meanwhile, Focal loss [44] and $L1$ loss are respectively used for the classification and 3D bounding box regression, the parameters of all mentioned modules are optimized together with the detector.
@@ III-B Pre-Fusion
The Pre-Fusion consists of two key modules: the VEM and the IEM. These two modules work together to enhance the semantic awareness of LiDAR features and the 3D spatial awareness of camera features, characterized by bidirectional complementarity, allowing for a more comprehensive 3D representation.
Figure 4: Image Enhancement Module (IEM). It first projects raw LiDAR points to the image coordinates. Then, we use depth completion to generate a dense depth feature map, whose size is the same as the camera features. Finally, the dense depth feature map is concatenated with the camera features, followed by a convolutional layer to reduce the feature dimension.
@@ III-B1 Voxel Enhancement Module
Despite having strong spatial localization capabilities, the LiDAR voxel features $F_{L}$ obtained from the LiDAR branch perform poorly on detailed semantic information. This drawback can propagate through the Unified Fusion process. To overcome this issue and provide more comprehensive features for the Unified Fusion, we design the VEM, as shown in Fig. 3.
@@
$\vspace{-1pt}F_{weighted}=\operatorname{Softmax}(L_{nearest})\times F_{nearest}\text{,}$ (1)
@@
$\vspace{-1pt}F_{SeL}=\operatorname{ReLU}(\operatorname{Linear}(F_{weighted}))+F_{L}\text{.}$ (2)
@@ III-B2 Image Enhancement Module
Camera features lack the ability to perceive accurate 3D spatial information, which is crucial for 3D object detection tasks. To enhance them from this view, we propose IEM, as shown in Fig. 4. We first project the point cloud onto the image plane to obtain a sparse depth map $D_{sparse}$. Inspired by [40], we then use depth completion [45] followed by feature extraction to obtain a dense depth feature map $D_{dense}\in\mathbb{R}^{H\times W\times C_{depth}}$. Finally, we concatenate the dense depth feature map with the camera features $F_{C}\in\mathbb{R}^{H\times W\times C_{2D}}$ and use a convolutional layer to fuse them, getting the spatial-aware camera features $F_{SpC}$:
$\vspace{-1pt}F_{SpC}=\operatorname{Conv}(\operatorname{Concat}(F_{C},D_{dense}))\text{.}$ (3)
Due to the geometric capability of the spatial-aware camera features, lifting them into 3D space ensures a more accurate process, providing a solid foundation for the subsequent Unified Fusion.
@@ III-C Unified Fusion
After obtaining the semantic-aware LiDAR voxel features and spatial-aware camera features, which have undergone specific bidirectional complementarity during Pre-Fusion, we further fuse them in the U-Fusion stage to build a semantic- and spatial-aware unified representation.
Previous work [1, 35] unifies the two modalities in the BEV space by view transformation. While this improved efficiency, it has an inherent limitation that prevents the fine-grained fusion which is vital for better performance. Therefore, after elevating the camera features to the voxel representation, we preserve its state, obtaining $\hat{F}_{SpC}$.